The speed of a stream is 3 mph. A boat travels 5 miles upstream in the same time it takes to travel 11 miles downstream. What is the speed of the boat in still water?


Any insight please.

Answer by mananth(16949)

boat speed x mph
current speed 3 mph
..
against current x - 3 mph
with current x + 3 mph

Distance upstream= 5
Distance downstream= 11
Time = same
5 /(x- 3 )= 11 /(x+ 3 )
5 *(x+ 3 )= 11 *(x- 3 )
5 x + 15 = 11 x -33
11 x -5 x = 33 + 15
6 x = 48
/ 6
x= 8 mph